OpenStudy (anonymous):

Solve 5(2x + 1) = 12x – 1. You must show all work.

OpenStudy (anonymous):

5(2x + 1) = 12x – 1 10x+5=12x-1 2x=6 x=3

OpenStudy (anonymous):

5(2x+1)=12x-1 10x+5=12x-1 6=2x x=3

OpenStudy (anonymous):

thank you